Calculate the magnitude of the faintest object that can be detected by an optical
telescope of diameter 5.0 m.
1
Expert's answer
2019-05-10T11:25:50-0400

Solution. The Scope's Star Magnitude Limit we can find using formula


"L_{mag}=2^m+5log(D_0)"

where Lmag is the magnitude limit of the scope; D0 is objective or main mirror diameter in mm. Therefore


"L_{mag}=2^m+5log(5000)=20.5^m"



Answer.

"L_{mag}=20.5^m"